OpenStudy (anonymous):

Find cos (2 theta angle) if sin (theta angle) =4/5

OpenStudy (anonymous):

Hint: $$ \cos 2 \theta = 1-2 \sin^2 \theta$$

OpenStudy (anonymous):

can you do the rest ?

OpenStudy (anonymous):

nah im aboslutely clueless about this question

OpenStudy (anonymous):

Hm. but I gave you the first clue .. it's just the matter of substitution now..

OpenStudy (anonymous):

Anyways the answer is $$ -\frac{7}{25}$$

OpenStudy (anonymous):

how you arrive there

OpenStudy (anonymous):

$$1-2\times (16/25)$$

OpenStudy (anonymous):

16 and 25, 1 and 2...where did all those numbers come from

OpenStudy (anonymous):

$$\cos 2 \theta = 1-2 \sin^2 \theta = 1-2\times (\frac 45)^2 = 1-2\times (16/25) = -\frac{7}{25}$$ anything else mam ?

OpenStudy (anonymous):

is the cos 2theta= 1-2sin^2theta a rule

OpenStudy (anonymous):

yes it's an identity.

OpenStudy (anonymous):

ok thanks for the help